Solve : dy/dx + y/x = y^2/x^2
1
Expert's answer
2020-09-16T19:22:19-0400

Solution:

To solve differential equation: "\frac{dy}{dx} + \frac y x=\frac{y^2}{x^2}"

"\frac{dy}{dx} + \frac y x=\frac{y^2}{x^2} ...(i)"

Put "y=vx ...(ii)"

On differentiating w.r.t "x"

"\frac{dy}{dx}=v+\frac{xdv}{dx} ...(iii)"

Put (ii) and (iii) in (i), we get,

"v+\frac{xdv}{dx}+\frac{vx}{x}=\frac{v^2x^2}{x^2}
\\ \Rightarrow v+\frac{xdv}{dx}+v=v^2
\\ \Rightarrow \frac{xdv}{dx}=v^2-2v
\\ \Rightarrow \frac{dv}{v^2-2v}=\frac{dx}{x}
\\ \Rightarrow \frac{dv}{(v-1)^2-1^2}=\frac{dx}{x}"

On integrating both sides,

"\frac 1 2 \ln |\frac{(v-1)-1}{(v-1)+1}|=\ln x+C
\\ \Rightarrow \frac 1 2 \ln |\frac{v-2}{v}|=\ln x+C
\\ \Rightarrow \ln |\frac{v-2}{v}|=2\ln x+2C"

"\\ \Rightarrow \ln |\frac{\frac y x-2}{\frac y x}|=2\ln x+2C" [Using (ii)]

"\\ \Rightarrow \ln |\frac{y-2x}{y}|=\ln x^2+2C 
\\ \Rightarrow \ln |\frac{y-2x}{y}|-\ln x^2=2C 
\\ \Rightarrow \ln |\frac{y-2x}{yx^2}|=2C"

"\\ \Rightarrow\frac{y-2x}{yx^2}=e^{2C}" [When "\ln y=x", then "y=e^x"]

"\\ \Rightarrow\frac{y-2x}{yx^2}=k" [where "k=e^{2C}", constant]

"\Rightarrow y-2x=kyx^2
\\ \Rightarrow y-kyx^2=2x
\\ \Rightarrow y(1-kx^2)=2x
\\ \Rightarrow y=\frac{2x}{1-kx^2}"

Answer:

The solution of given differential equation is "y=\frac{2x}{1-kx^2}" , where "k" is a constant.